This is a word problem. One number is 12 more than another. The sum of the smaller number and twice the larger number is 39. Find the larger number.
I do not even know where to begin on this problem. Please Help!!!!!

Answer by stanbon(75887)

This is a word problem. One number is 12 more than another. The sum of the smaller number and twice the larger number is 39. Find the larger number.
-----------------
Let the two numbers be "x" and "x+12".
EQUATIONS:
x+2(x+12)=39
2x+2x+24=39
4x=27
x=27/4 (this is the smaller number)
x+12=27/4 + 48/4 = 75/4 (this is the larger number)
Cheers,
Stan H.